OpenStudy (anonymous):

what number makes this equation true? v^2+10v+16=(v+8)(v+?)

OpenStudy (texaschic101):

what number added to 8 gives you 10 and multiplied by 8 gives you 16 ?

OpenStudy (gudden):

@mena_loves_you . divide the Equation with (v+8) as the first step..

OpenStudy (anonymous):

2 right? Sorry my computer is being a pain in the A$$

OpenStudy (gudden):

v^2 + 2v + 8v + 16 = (v+8) ( v+?) v(v+2) + 8 (v + 2) = (v+8) ( v + ? ) (v+2) ( v + 8 ) = ( v+8) (v+?) dividing (v+8) on both the sides we get (v+2) = ( v + ? ) i.e. v + 2 = v + ? subtracting v from both sides we get 2 = ? hence missing number = 2

OpenStudy (anonymous):

Omg thank u so much (:

OpenStudy (gudden):

ur welcome...